You just acquired a mortgage in the amount of $249,500 at 5.75 percent interest, compounded monthly. Equal payments are to be made at the end of each month for thirty years. How much of the first loan payment is interest? (Assume each month is equal to 1/12 of a year.)
Answer #1

Monthly payment=PMT(5.75%/12,12*30,-249500,0)=$1,456.01

Interest payment=249500*5.75%/12=$1,195.52

% of interest payment=1195.52/1456.01=82.109%